OpenStudy (anonymous):

Finding area of trapezoid. With no hight. Base 9ft and 12 ft and leg of 9 and 8 ft

Directrix (directrix):

@Jennyvcr 9^2 = h^2 + x^2 8^2 = h^2 + (3 - x) ^2 Solve those simultaneously. The object is to get h so that you can find the area of the trapezoid.